OpenStudy (anonymous):

If y varies inversely as x and y = 10 when x = 5, then the constant of variation would be: a) 100 b)5 c)2 d) 50

OpenStudy (anonymous):

\[y = c(1/x)\], if y = 10 when x = 5, you would have 10 = c(1/5), c = 50.